Cars A and B, each moving at 70 km/h are approaching one another on a straight highway. After how many seconds will they meet if they were initially 1 km apart?
Answer by mananth(16946)

d=1km
rate = 70 km/h of both
moving towards each other
so effective rate = 140 km/h
t=d/r
t=1/140 hour
t(sec) = 1/140 * 3600
25.71 seconds